# Find Any Part in 3D

### v1

@@ 4 A General Open-World 3D Part Benchmark
3D part segmentation has largely been shaped by benchmarks like ShapeNet-Part [29] and PartNet-E [14, 28], which annotate CAD models from a small number of categories, such as chairs and tables, with a fixed set of parts. Additionally, these benchmarks position objects in canonical orientations (e.g., all chairs face right), biasing the trained methods to handle only canonically oriented objects. We introduce a new benchmark featuring a diverse range of objects, shapes, parts, and poses to advance 3D object part segmentation toward more general objects and parts, without assumptions about object orientation. Our hope for this benchmark is to advance real-world applicability of 3D models in more variable, “in-the-wild” scenarios.
We source 3D objects from the large Objaverse dataset [3] with a variety of object types and shapes. The benchmark is divided into two sets described below.
@@
Objaverse-General is a dataset of 100 objects from 100 diverse object categories, such as gondola, slide, lamppost, easel, penguin. This set is labeled with a total of 350 parts, annotated by humans to guarantee high quality annotations. These objects are in random orientations.
Figure 4: Qualitative results on the Objaverse-General benchmark. Find3D demonstrates strong performance on diverse objects, both for seen and unseen categories. PointCLIPV2, trained on ShapeNet-Part, generalizes poorly to novel object categories. PartSLIP++, a detection-based method, predicts sparsely because many parts are not successfully detected. OpenMask3D cannot go below the object-level granularity, and usually assigns the whole object to one part.
Dataset # Objects # Categories # Parts
Objaverse-General 100 100 350
Objaverse-ShapeNetPart 32 16 90
Table 1: Statistics of the two benchmarks we propose for general, open-world 3D part segmentation.
mIoU (%) Objaverse-General ShapeNet-Part PartNet-E
Time Seen Categories Unseen Categories Overall Canonical Rotated Objaverse-ShapeNetPart Canonical Rotated
Open-W Find3D (ours) 0.9s 33.78 26.21 29.99 28.39 29.64 42.15 16.86 17.62
PointCLIPV2 5.4s $9.81$ $10.27$ $10.04$ 16.91 16.88 $15.14$ $11.28$ $10.32$
PartSLIP++ 174.3s $2.69$ $0.57$ $1.63$ $1.43$ $0.93$ $1.54$ 5.12 3.87
OpenMask3D 296.5s $11.81$ $7.01$ $9.41$ $8.94$ $6.75$ $15.87$ $12.54$ $11.93$
Closed-W
PointNext 80.44 32.18 $28.70$
PartDistill 63.9** 39.9**
Table 2: Performance comparison on Objaverse-General, ShapeNet-Part, and PartNet-E. Find3D is trained on Objaverse, PointCLIPV2 and PointNext are trained on ShapeNet-Part, and PartSLIP++ is trained on PartNet-E (shaded). Besides PartDistill which is category-specific, all methods are evaluated zero-shot on other datasets. Among the open-world methods, Find3D performs best on all datasets, achieving up to 3x higher mIoU and 6-300x faster inference time. “-” denotes that closed-world methods cannot be evaluated on datasets containing unseen categories. We report class-average mIoU evaluated with a fixed prompt (“{part} of a {object}”). Additional evaluations with varying query prompts can be found in the appendix. ** means metric not reproducible and directly taken from papers.
@@
Objaverse-ShapeNetPart contains 32 objects from the same 16 object types as in ShapeNet-Part [29]. Unlike ShapeNet-Part, the objects come in random orientations. This benchmark is used to evaluate robustness with regard to distribution shift for models trained on ShapeNet-Part.
Tab. 1 provides statistics of our proposed benchmarks. To comprehensively evaluate generalization of our model, we randomly select 50 out of the 100 Objaverse-General categories. We hold out all objects of these categories, as well as visually similar categories during training. We evaluate our model both on the 50 held-out categories (Unseen-Categories) and the other 50 categories (Seen-Categories). All 100 instances in the benchmark are held out during training to ensure no data leakage during evaluation.
@@ 5 Experiments
Our open-world 3D part segmentation experiments show:
  • Find3D obtains strong performance on generic object categories both for seen and unseen categories, achieving more than 3x mIoU of the second best method. Find3D exhibits strong out-of-distribution generalization, whereas baseline methods perform poorly on datasets they are not trained on, as shown qualitatively in Fig. 4 and quantitatively in Tab. 2.
  • Find3D is robust to variations such as query text prompt rephrasing and object rotation, whereas baselines are sensitive to these changes. This is shown qualitatively in Fig. 5 and quantitatively in Fig. 6.
  • Find3D can segment single-image reconstructions from in-the-wild images, such as iPhone photos and DALLE2 generations, as shown in Fig. 1 (panel c).
  • Find3D is the most efficient method – 6x to over 300x faster than open-vocabulary baselines, as shown in Tab. 2.
@@ 5.1 Experimental Settings
Benchmarks. In addition to our proposed benchmark (Sec. 4), we also evaluate on two commonly used datasets for 3D part segmentation: ShapeNet-Part [29] and PartNet-E [14]. ShapeNet-Part contains 16 object categories with canonical poses. PartNet-E combines PartNet [14] and PartNet-Mobility [28]. Its test set contains 45 household object classes. For both datasets, we evaluate on their test set both in the canonical pose and in a randomly rotated (around all axes) pose, forming four additional benchmarks: ShapeNetPart-Canonical, ShapeNetPart-Rotated, PartNetE-Canonical, and PartNetE-Rotated.
@@
Competing Methods. We compare to state-of-the-art open- and closed-world methods for 3D segmentation.
Open-world Baselines: We compare to three open-world methods. PointCLIPV2 [35] is an open-world 2D-to-3D pipeline involving multiple invocations of CLIP [18] using top-k ($k\geq 1400$) prompts selected on the test set. The pipeline is tuned on ShapeNet-Part. PartSLIP++ [34] is detection-based pipeline involving invocations of GLIP [9] and an algorithm for finding superpoints to improve segmentation. PartSLIP++ is trained on PartNet-E. We evaluate its zero-shot checkpoint for fairness of comparison. OpenMask3D [22] is an open-vocabulary, 2D-to-3D pipeline trained on holistic scenes. PointCLIPV2 and OpenMask3D are dense methods that assign a label to every point. On benchmarks that contain unlabeled points, we provide the text query “other” as a option for no label.
Closed-world Baselines: We compare to two closed-world methods: PointNeXt [17] is a state-of-the-art closed-world point cloud segmentation model trained on ShapeNet-Part. Due to its closed vocabulary, it cannot be evaluated on all datasets. PartDistill [23] is a category-specific 2D-to-3D distillation method, which is open-world in the 2D stage but closed-world in the 3D stage. It cannot be evaluated on unseen object categories due to the category-specific nature of distillation. The code and data for this method are not fully released (only two categories are released). Since we cannot reproduce the approach, we show numbers claimed in the paper.
Because PartSLIP++ and OpenMask3D are slow (up to 5 minutes per object), they are infeasible to evaluate on the full test sets. For example, evaluating OpenMask3D on PartNet-E test set would take 628 hours. To facilitate complete and fair evaluations of all methods, we create smaller subsets of 160 objects (10 objects/category $\times$ 16 categories) for ShapeNet-Part and 225 objects (5 objects/category $\times$ 45 categories) for PartNet-E. For methods that are efficient to evaluate, we additionally report performance on the full test sets. We observe similar results and rankings of methods as on the subsets and full sets (when applicable). Complete results and the subset indices are provided in the appendix.
@@ 5.2 Experimental Results
[…]
@@
Fig. 4 provides a qualitative comparison on Objaverse-General. Our method consistently outputs reasonable segmentations, while other methods struggle. PartSLIP++ is detection-based, and predicts “no label” if nothing is detected. It is trained on PartNet-E with sparse part annotations, and thus tends to predict sparsely. OpenMask3D struggles with the part-level granularity, and it usually only picks one part, or at most two parts, to represent the whole object. We do note that our method tends to generate dense predictions, and only rarely predicts “no label”.
@@ 5.2.1 Quantifying and Comparing Robustness
We evaluate the robustness of our method by varying the query text prompt, the object orientation, as well as the data domain, i.e. the data source of similarly-looking objects. Fig. 5 provides qualitative examples.
@@
Figure 5: Robustness comparison of PointCLIPV2 and Find3D. We evaluate both methods on a ShapeNet-Part earphone with canonical and rotated orientations, and a visually similar earphone from Objaverse-ShapeNetPart. Top-k prompt reproduces evaluation in the PointCLIPV2 paper. PointCLIPV2’s performance drops up to $68\%$ with variations in evaluation configurations, whereas our method stays consistent.
Figure 6: Robustness evaluation of PointCLIPV2 and Find3D on all ShapeNet-Part categories. We vary the query prompt, object orientation and data source (same object categories from ShapeNet-Part vs. Objaverse-ShapeNetPart). PointCLIPV2’s performance drops up to $64\%$, while our method remains robust.
@@
Robustness to query prompt. PointCLIPV2’s reported metric comes from an extensive top-k prompt search on the test set: for each object part, they try 700 LLM-generated prompts. They iterate 2 times over all parts on the test set to determine the best prompt. For fairness of comparison, we also perform this top-k search for our method. Additionally, we evaluate on two prompts that are more likely to be used in a real application: the “part of a object” prompt (such as “leg of a chair”), or the “part name” prompt (“leg”). As shown in Fig. 6, with a change of prompt from top-k to “part of a object”, PointCLIPV2’s performance drops from 48.472 to 17.417, i.e., a $64\%$ drop of performance, whereas our method only drops $31\%$, staying much more robust.
@@
Robustness to object orientation. To evaluate robustness to object orientation, we apply a random rotation by sampling three angles from $-\pi$ to $\pi$ and applying rotations along each of the X, Y, Z axis sequentially. As shown in Fig. 6, PointCLIPV2’s performance drops $46\%$ whereas our method does not drop but even increases $3\%$.
@@
Robustness to domain. We constructed Objaverse-ShapeNetPart, a benchmark with objects from the same categories as ShapeNet-Part, but sourced from Objaverse assets. With this domain shift, PointCLIPV2’s performance drops from 48.472 to 21.177 by $56\%$, whereas our method stays robust with a $20\%$ increase. We provide more comprehensive evaluations with varying configurations on the other datasets in the appendix, which show similar trends.
5.2.2 In-the-Wild Results
While varying evaluation settings on benchmark datasets allows for a systematic study of robustness, the ultimate test of robustness is whether a method can work in-the-wild. To this end, we demonstrate segmentation results on 3D reconstructions from iPhone photos or DALLE-2 generations. We also show the capability to handle flexible text queries in terms of granularity and description style.
Segmenting 3D reconstructions of in-the-wild images. Point clouds are relatively easy to obtain in the wild – for example, with single-image reconstruction techniques like MCC [25]. Find3D can work on noisy point clouds reconstructed from iPhone photos or AI-generated images. In Fig. 1, Find3D successfully segments a push car reconstructed from an iPhone photo, and a reconstructed point cloud from the DALLE2-generated image of “a capybara wearing a hat”. These results demonstrate that Find3D has great potential in real applications – it works with everyday objects that are very different from the synthetic training data, and works with 3D reconstructions that are noisy.

### v2

@@ 4 A General Open-World 3D Part Benchmark
Figure 4: Our benchmark. (a) Examples of Objaverse-General and ShapeNetPart-V2. Objaverse-General contains diverse objects and parts, and ShapeNetPart-V2 is sourced to look similar to ShapeNet-Part to test various methods’ generalization capability. (b) Object category breakdown of Objaverse-General, which covers 9 categories from tools to buildings. (c) Comparison with existing benchmarks. We have $\mathbf{5\times}$ more unique part types, $\mathbf{4.4\times}$ more total annotated parts, and $\mathbf{2.9\times}$ more object categories.
Existing 3D part segmentation benchmarks only contain a small number of categories with a fixed set of parts, and are limited to narrow domains. ShapeNet-Part [27] contains 16 object categories with 41 unique part types, and the domain is limited to CAD models with canonical orientations (e.g., all chairs face right). PartNet-E [12, 24] contains 45 categories with 40 unique part types (e.g., “button” is a common part across categories), and is restricted to simple home objects such as bottles and doors. As shown in  Fig. 4, we introduce a new human-annotated benchmark featuring a diverse range of objects, shapes, parts, and poses. We source our data from Objaverse [2]. Our benchmark contains 132 object categories, 450 total parts and 207 unique part types, over $\mathbf{5\times}$ that of existing benchmarks, as shown in Fig. 4. We hope this benchmark can advance 3D part segmentation towards more variable, “in-the-wild” scenarios. The benchmark is divided into two sets:
@@
Objaverse-General contains 100 objects with 350 parts from 100 diverse object categories, such as gondola, slide, lamppost, easel, penguin. These objects are in random orientations. We hold out 50 out of the 100 categories from training, in order to evaluate out-of-distribution generalization to novel object types. The holdout categories are termed Unseen-Categories in Table 2.
@@
ShapeNetPart-V2 contains 32 objects from the same 16 object categories in ShapeNet-Part [27]. Inspired by ImageNetV2 [16], we create this benchmark to evaluate generalization for models trained on ShapeNet-Part.
@@ 5 Experiments
Method Time Open-world Cross-category Part-level Feed-forward
Find3D 000.9s
PointCLIPV2 005.4s
PartSLIP++ 174.3s
OpenMask3D 296.5s
PartDistill 000.7s (+348s)
PointNext 001.4s
Table 1: Properties and inference time of Find3D and baselines. Find3D is the only method that is open-world, cross-category, part-level and feed-forward. By “feed-forward”, we mean a method that performs direct 3D inference, without relying on a pipeline of multiview rendering, multiple 2D model inferences, and backprojection. Our method is $\mathbf{6\times}$ to $\mathbf{300\times}$ faster than other open-world models, and on par with closed-world models.
$\ast$ PartSLIP++ finetunes a model per category. $\dagger$ PartDistill performs distillation for each new category (348s) and then does inference (0.7s). It is not open-world as the part names need to be defined prior to distillation. It only releases source code for two categories, and our reported speed is averaged across them.
Figure 5: Qualitative results. Left: Find3D performs strongly on Objaverse-General while baseline methods struggle. Right: more examples both from Objaverse-General and PartObjaverse-Tiny, including out-of-distribution objects such as magical animals and complex anime-style characters. Find3D works on diverse object categories with up to 9 parts. It also generalizes to “in-the-wild” iPhone photos (converted to point clouds via off-the-shelf image-to-3D method, as shown at bottom right.
As summarized in Table 1, Find3D is the first method that is simultaneously, open-world, cross-category, part-level and feed-forward. Find3D not only shows strong zero-shot generalization, but also outperforms existing methods on their own domain. Our experiments show:
  • Find3D achieves strong performance on diverse objects, with $\mathbf{260\%}$ improvement in mIoU from existing methods. Find3D exhibits strong out-of-distribution generalization, whereas baseline methods perform poorly on datasets they are not trained on, as shown qualitatively in Fig. 5 and quantitatively in Tab. 2, Tab. 3, Tab. 4.
  • Find3D is robust to variations such as query prompt rephrasing, object rotation, and domain shift, whereas baselines are sensitive to these changes. This is shown in Fig. 7, Tab. 3, Tab. 4.
  • Find3D is the most efficient open-world method with 6x to 300x speed improvement, as shown in Tab. 1.
@@ 5.1 Experimental Settings
Benchmarks. In addition to our proposed benchmark (Sec. 4), we also evaluate on two commonly used datasets for 3D part segmentation: ShapeNet-Part [27] (16 object categories) and PartNet-E [12] (45 object categories). For both datasets, we evaluate on their test set both in the canonical pose and in a randomly rotated (around all axes) pose, which correspond to the Canonical and Rotated columns in Tab. 2, Tab. 3.
mIoU (%) Objaverse-General ShapeNet-Part
Seen Categories Unseen Categories Canonical Rotated ShapeNetPart-V2
{part} of a {object} {part} {part} of a {object} {part} {part} of a {object} {part} {part} of a {object} {part} {part} of a {object} {part}
Find3D (ours) 33.78 34.10 26.21 27.41 28.39 24.09 29.64 23.71 42.15 30.02
PointCLIPV2 $9.81$ $11.27$ $10.27$ $11.09$ 16.91 20.22 16.88 18.19 $15.14$ $17.11$
PartSLIP++ $2.69$ $15.03$ $0.57$ $10.43$ $1.43$ $6.46$ $0.94$ $6.03$ $1.54$ $11.62$
OpenMask3D $11.81$ $11.93$ $7.01$ $10.31$ $8.94$ $10.37$ $6.75$ $14.56$ $15.87$ $13.77$
Table 2: Performance comparison with open-world methods on Objaverse-General and ShapeNet-Part. Shaded cells mean the method is trained on the same dataset (expected higher than white cells); white cells mean zero-shot evaluation. Find3D performs best on Objaverse-General, with $\mathbf{260\%}$ improvement in mIoU on unseen categories where all methods are evaluated zero-shot. On ShapeNet-Part, Find3D’s zero-shot performance even exceeds PointCLIPV2 which is trained on the this dataset. We show results evaluated with 2 common query prompts: “{part} of a {object}” and “{part}” for all methods.
@@
Competing Methods
Open-world Baselines: PointCLIPV2 [33] is an open-world 2D-to-3D pipeline involving multiple invocations of CLIP [15]. It uses top-k prompts ($k=1400\times n_{\text{parts}}$ per object) selected on the test set of ShapeNet-Part. PartSLIP++ [32] is a detection-based pipeline involving invocations of GLIP [7] and a custom algorithm for finding superpoints. It finetunes a separate model for each category in PartNet-E. We evaluate its zero-shot checkpoint for fairness of comparison. OpenMask3D [20] is an open-vocabulary, 2D-to-3D pipeline trained on scenes.
PointCLIPV2 and OpenMask3D are dense methods that assign a label to every point. We provide the text query “other” as an option for no label on benchmarks that contain unlabeled points.
Closed-world Baselines: PointNeXt [14] is a state-of-the-art closed-world point cloud segmentation model trained on ShapeNet-Part. Due to its closed vocabulary, it cannot be evaluated on other datasets. PartDistill [21] is a category-specific 2D-to-3D distillation method, which is open-world prior to distillation but closed-world at inference time. It cannot be evaluated on unseen object categories due to the category-specific nature of distillation. The code and data for this method are not fully released (only two categories are released). Since we cannot reproduce the approach, we show numbers claimed in the paper.
Because PartSLIP++ and OpenMask3D are slow (up to 5 minutes per object), they are infeasible to evaluate on the full test sets (evaluating OpenMask3D on PartNet-E test set would take 628 hours). For fair evaluations of all methods, we create smaller subsets of 160 objects (10 objects/category $\times$ 16 categories) for ShapeNet-Part and 225 objects (5 objects/category $\times$ 45 categories) for PartNet-E. For methods that are efficient to evaluate, we additionally report performance on the full test sets in the appendix. We observe the same rankings and similar results on the subsets and full sets.
@@ 5.2 Experimental Results
Results on Objaverse-General and ShapeNet-Part. Tab. 2 reports the mIoU of Find3D and open-world baselines on Objaverse-General and ShapeNet-Part. Find3D shows the strongest performance, with $\mathbf{260\%}$ improvement in mIoU compared to the best baseline, PointCLIPV2, when both are evaluated zero-shot out of distribution (Objaverse-General– Unseen Categories). Additionally, even when evaluated zero-shot, Find3D outperforms PointCLIPV2 on ShapeNet-Part, the dataset it is trained on.
@@
Qualitative results. As seen in Fig. 5, Find3D consistently outputs reasonable segmentations, while other methods struggle. PartSLIP++ is trained on PartNet-E with sparse part annotations, and thus tends to output “no label” overly often. OpenMask3D struggles with the part-level granularity, and it usually only picks one part, or at most two parts, to represent the whole object. We additionally show examples in Objaverse-General and PartObjaverse-Tiny [26] for our method on the right. Find3D not only works with diverse objects and parts, but also generalizes to real-world iPhone photos (converted to point clouds with Trellis [25], an off-the-shelf image-to-3D model), despite only being trained on synthetic assets.
Results on PartNet-E.
mIoU(%) Canonical Orientation Rotated
{part} of a {object} {part} {part} of a {object} {part}
Find3D (ours) 16.86 16.38 17.62 17.16
PartSLIP++ 05.12 32.71 03.87 23.03
PointCLIPV2 11.28 09.70 10.32 10.22
OpenMask3D 12.54 11.24 11.93 11.67
Table 3: Comparison of open-world methods on PartNet-E. Shaded cells mean the method is trained on the same dataset (expected higher than white cells); white cells mean zero-shot evaluation. We evaluate with 2 prompt formats: “{part} of a {object}” and “{part}”. PartSLIP++ achieves good performance with the “{part}” prompt, but its performance drops $\mathbf{84\%}$ when we vary the query prompt. This dataset is challenging for our method due to the sparsity of labels and the presence of small parts that are not geometrically or colorfully prominent (e.g., buttons on a surface with the same color). Nevertheless, our method is more robust to rotation and prompt variation, and clearly outperforms the other baselines that are also evaluated zero-shot.
Tab. 3 compares open-world methods on PartNet-E. PartSLIP++ is trained on this dataset while all other methods, including Find3D, are evaluated zero-shot, so the results in this table favor PartSLIP++. We evaluate on two common prompts: “part of a object” (such as “leg of a chair”), and “part name” (“leg”). We see that PartSLIP++’s performance decreases greatly when we vary the query prompt, up to a $83\%$ drop. Our method, evaluated zero-shot, is more robust and outperforms PartSLIP++ with the “part of a object” prompt. It also outperforms other zero-shot baselines under all evaluation configurations.
Efficiency. As shown in  Tab. 1, Find3D only takes 0.9 seconds for inference, which is 6$\times$ to 300$\times$ faster than open-world baselines and on par with closed-world models. Inference time is the average per-object inference time on the PartNet-E subset evaluated on an A100.
Comparing with closed-world methods on ShapeNet-Part and ShapeNetPart-V2. Tab. 4 compares Find3D zero-shot with closed-world methods that are trained on ShapeNet-Part, which greatly favors the closed-world methods. PointNeXt is the leading closed-world method for this dataset, and PartDistill trains one model for each object category of ShapeNet-Part. We additionally evaluate the methods’ generalization capability on our ShapeNetPart-V2 benchmark, similar to ImageNetV2 [16]. We see a $\mathbf{64\%}$ drop of PointNeXt. Even though PointNeXt is still in-domain and Find3D is evaluated out-of-distribution, Find3D shows a $\mathbf{1.5\times}$ advantage. PartDistill is not reproducible and thus cannot be evaluated on ShapeNetPart-V2.
mIoU (%) Trained on ShapeNet-Part ShapeNetPart-V2
Find3D Our data engine $28.39$ $42.15$
PointNeXt ShapeNet-Part 80.44 $28.70$
PartDistill ShapeNet-Part 63.9 N/A
Table 4: Performance comparison with closed-world methods. Shaded cells mean the method is trained on the same dataset (expected higher than white cells); white cells mean zero-shot evaluation. PointNeXt, a state-of-the-art closed-world model, is trained on ShapeNet-Part, but its performance drops significantly on ShapeNetPart-V2. Our approach, which is trained on a domain different from either ShapeNet-Part and ShapeNetPart-V2, demonstrates a stronger out-of-domain performance (1.5×1.5$\times$1.5 × better on ShapeNetPart-V2, +13.2%). $\dagger$ PartDistill trains a model per-category on ShapeNet-Part. It does not release training source-code or checkpoints (apart from two categories), thus cannot be evaluated on ShapeNetPart-V2.
5.2.1 Scaling Analysis
Figure 6: Data scaling which shows that training on more object categories provides clear improvement in zero-shot mIoU. The evaluation is done on Objaverse-General Unseen Categories.
Data scaling is critical, as shown by the scaling analysis in Fig. 6. This finding highlights the importance of our data engine approach, which enables scaling in 3D. We vary training object categories (x-axis) ranging from 16 categories (ShapeNet-Part dataset size), 45 categories (PartNet-E dataset size), all the way to 761 (our setting). We report zero-shot mIoU on Objaverse-Unseen Categories (y-axis). We observe a strong scaling trend which is consistent with findings in many other data domains.
@@ 5.2.2 Quantifying and Comparing Robustness
Fig. 7 evaluates the robustness of our method under different query text prompt, object orientation, and data domain, i.e. the data source of similarly-looking objects.
@@
Figure 7: (a) Qualitative comparison of PointCLIPV2 and Find3D on a ShapeNet-Part earphone (canonical and rotated) and a visually similar earphone from ShapeNetPart-V2. Top-k prompt reproduces evaluation in the PointCLIPV2 paper. PointCLIPV2’s performance drops up to $68\%$, whereas our method stays consistent. (b) Comparison over all ShapeNet-Part categories. PointCLIPV2’s performance drops $46\%$ to $64\%$ with varying conditions, while our method remains robust.
@@
Robustness to query prompt. PointCLIPV2 performs an extensive top-k prompt search on the test set: they iteratively optimize the prompt for each part (iteratively searching over 700 prompts per part and looping over all parts twice) and select the best prompt, i.e. iterate over $k=1400\times n_{\text{part}}$ prompts and pick the one with the best test performance. For a fair comparison, we perform the same top-k search for our method. We also evaluate on two common prompts: “part of a object” (such as “leg of a chair”), and “part name” (“leg”). As shown in Fig. 7, with a change of prompt from top-k to “part of a object”, PointCLIPV2’s performance drops from 48.47 to 17.42 ($64\%$ decrease), whereas our method exhibits more robust performance.
@@
Robustness to object orientation. We apply a random rotation by sampling three angles from $-\pi$ to $\pi$ and applying rotations along each of the X, Y, Z axis sequentially. PointCLIPV2’s performance drops $46\%$ whereas our method does not drop but even increases $3\%$.
@@
Robustness to domain. We constructed ShapeNetPart-V2, a benchmark with objects from the same categories as ShapeNet-Part, but sourced from Objaverse assets. With this domain shift, PointCLIPV2’s performance drops from 48.47 to 21.18 by $56\%$, whereas our method stays robust with a $20\%$ increase.
Comparison with other methods on other datasets in  Tab. 2 and  Tab. 3 show similar trends.
5.2.3 Flexibility of text queries